OpenStudy (anonymous):

square root of 3 times the square root of 25. simpify? please help.

OpenStudy (binary3i):

5root5

OpenStudy (anonymous):

\[\Large \sqrt 3 \cdot \sqrt {25} = 5\sqrt 3 \]

OpenStudy (anonymous):

\[\sqrt{3} \sqrt{25} = 5\sqrt{3}\]

OpenStudy (anonymous):

Beat me to it

OpenStudy (angela210793):

wht's the right one for God's sake -_-

OpenStudy (binary3i):

im mistakn

OpenStudy (anonymous):

I'm assuming that's what she meant; angela is correct but it's a bit ambiguous with "the square root of blah blah blah" rather than using brackets etc ;)

OpenStudy (anonymous):

It is sqrt(15)

OpenStudy (anonymous):

\[\sqrt{15}\]

OpenStudy (anonymous):

I've never been able to find a solution for "blah blah blah"

OpenStudy (angela210793):

Ahhhhha...i got it....y do i always have 2 make things more complicated than they really are -_- :(((((

OpenStudy (anonymous):

The answer is sqrt(15) for pete's sake

OpenStudy (anonymous):